The weight of a rock is measured to be 0.055 lbs. After placing the rock in a graduated cylinder containing 13.99 mL of water, the new measurement is 29.51 mL. Calculate the density of the rock in g/mL.


1
Expert's answer
2021-09-10T00:42:16-0400

1 lbs =

453.592 g

= 0.055 × 453.592 = 24.95 g

V = 29.51-13.99 = 15.52ml

density = 24.95/15.52 = 1.61g/ml